OpenStudy (anonymous):

Differential Equation help. Newtons’ law of cooling tells us that the rate at which a body cools is proportional to the difference between the cooling body and the ambient temperature, To . This gives the equation dT/dt=k(T-To) As T is decreasing k<0. If T=60 at t=0 T=50 at t=20 and T=45 when t=40 find the temperature when t=30.

OpenStudy (anonymous):

dont you need some kind of t on left side? I think?

OpenStudy (anonymous):

You need to separate the variables. Int(dT/(T-T0))=Int(k)dt and solve for the constant to find the equation

OpenStudy (anonymous):

t-Intat(1/k(_a-To), _a = T(t))+_C1 = 0

OpenStudy (anonymous):

hmm i think this question is worded weirdly, do you think they meant that To is t?

OpenStudy (anonymous):

T0 probably means the first number in the sequence.

OpenStudy (anonymous):

so after doing the integrals i got ln(T-To)=kt+c I can sub in T and t but not sure what to put for To?

OpenStudy (anonymous):

and further simplification i get T=Ae^(kt)+To

OpenStudy (anonymous):

oh i think i can get it from now given that i have three temps to play with i can find A and To

OpenStudy (anonymous):

Yes, I would substitute in. My guess is that T0 is the initial position or whatever T is defined as.

OpenStudy (anonymous):

equation for T is apparently, T=40+20e^(tln2/20)....so hopefully i can get there. Does look like To=t here....just not sure how that makes sense...

OpenStudy (anonymous):

no no To is not t.... dont know what i was thinking!

OpenStudy (anonymous):

I end up with this system of equations. 60=A+To 50=Ae^(20k)+To 45=Ae^40k)+To but im stuck as to how to then get k, To and A

OpenStudy (anonymous):

hang on might be on to it...sorry....

OpenStudy (anonymous):

ok got it. Thanks heaps for getting me started!! So for anyone interested in the conclusion, equations are rearranged so that To=60-A, and A=(-10/(e^(20k)-1) and k is solved by solving for 45 when the rhs has only the k variable. This gives A=20, k=-ln(2)/20 and To=40 which in T, is T=40+20e^(-tln(2)/20). And finally, when t=30, T=40+5sqrt2 Thanks again!